11. (7 pts.) Compute the Jacobian for the transformation x = ue and y=ue. Bonus: Find the (u, v) coordinates for the point

  1. Compute the Jacobian for the transformation and.

Bonus: Find the coordinates for the point in the xy-Plane.
Answer #1

x=uev y=ueu ax dx au Jacobian, d(x,y) 11 d (Usv) 2 y ду av du uer 11 en - uer sex é V-v ee as =-u-u o = e =-2u | Answer The p